# Trans-Glasso: A Transfer Learning Approach to Precision Matrix Estimation

### v1

@@ Abstract
Precision matrix estimation is essential in various fields, yet it is challenging when samples for the target study are limited. Transfer learning can enhance estimation accuracy by leveraging data from related source studies. We propose Trans-Glasso, a two-step transfer learning method for precision matrix estimation. First, we obtain initial estimators using a multi-task learning objective that captures shared and unique features across studies. Then, we refine these estimators through differential network estimation to adjust for structural differences between the target and source precision matrices. Under the assumption that most entries of the target precision matrix are shared with source matrices, we derive non-asymptotic error bounds and show that Trans-Glasso achieves minimax optimality under certain conditions. Extensive simulations demonstrate Trans-Glasso ’s superior performance compared to baseline methods, particularly in small-sample settings. We further validate Trans-Glasso in applications to gene networks across brain tissues and protein networks for various cancer subtypes, showcasing its effectiveness in biological contexts. Additionally, we derive the minimax optimal rate for differential network estimation, representing the first such guarantee in this area.
@@ 1 Introduction
Estimating the precision matrix, i.e., the inverse covariance matrix, is a fundamental task in statistical analysis and has broad applications, including in portfolio optimization, speech recognition, and genomics [Best and Grauer, 1992, Lauritzen, 1996, Yuan and Lin, 2007, Saon and Chien, 2011]. The precision matrix is closely tied to Gaussian graphical models: estimating the support of the precision matrix corresponds to uncovering the network structure of conditional dependencies between multivariate normal variables [Lauritzen, 1996]. However, estimating a precision matrix accurately is often challenging when the sample size is small compared to the dimension—a typical scenario in high-dimensional settings.
In many applications, sample sizes are constrained for the target study of interest, yet data from related studies may be available. Transfer learning [Pan and Yang, 2009] provides a promising approach in these scenarios by leveraging information from related source studies to improve estimation accuracy in the target study. For example, in gene expression studies across different tissue types, sample sizes may be small for specific tissues, but data from related tissues can help improve estimates [Li et al., 2023a]. Similarly, protein network studies for different cancer subtypes can benefit from transfer learning, as leveraging data from related subtypes can enhance estimation for a particular subtype with limited data [Peterson et al., 2015].
A critical aspect of transfer learning is establishing similarity between the target and source tasks. Here, we assume that most entries of the target precision matrix are shared with those of the source matrices, with only a few differences. Based on this assumption, we propose Trans-Glasso, a novel two-step transfer learning method for precision matrix estimation. First, we obtain initial estimators through a multi-task learning objective that captures shared and unique dependencies across datasets. Second, we refine these estimators using differential network estimation to adjust for differences between the target and source matrices [Zhao et al., 2014, Yuan et al., 2017].
We provide a theoretical analysis of Trans-Glasso, deriving non-asymptotic error bounds and establishing that the method achieves minimax optimality in a wide range of parameter regimes. Through extensive simulations, we demonstrate that Trans-Glasso outperforms several baseline methods, particularly in scenarios where the target sample size is small. We also apply Trans-Glasso to gene networks across brain tissues and protein networks for various cancer subtypes, showing its practical effectiveness in biological applications. Additionally, as a byproduct of our analysis, we derive the minimax optimal rate for differential network estimation, to our knowledge, the first of its kind.
@@ 1.1 Related Work
Estimation of sparse precision matrices in a single study is well studied. Common methods include penalized M-estimator [Yuan and Lin, 2007, Friedman et al., 2008, Rothman et al., 2008, Lam and Fan, 2009, Ravikumar et al., 2011] and constrained $L_{1}$ minimization [Cai et al., 2011, Ren et al., 2015, Cai et al., 2016b]. There is also extensive literature on multi-task precision matrix estimation, which estimates multiple related but nonidentical precision matrices from multiple studies [Guo et al., 2011, Danaher et al., 2014, Zhu et al., 2014, Mohan et al., 2014, Lee and Liu, 2015, Cai et al., 2016a, Ma and Michailidis, 2016, Saegusa and Shojaie, 2016]. See Tsai et al. [2022] for a survey. While related to transfer learning, multi-task learning aims to estimate parameters of all studies, whereas transfer learning only focuses on the target study.
@@
Transfer learning has a long history [Pan and Yang, 2009] and has been applied in various contexts [Turki et al., 2017, Hajiramezanali et al., 2018, Bastani, 2021]. Recently, interest in transfer learning for statistical problems has grown. Li et al. [2022], He et al. [2024] studied high-dimensional linear regression. The fused regularizer in He et al. [2024] is similar to our multi-task objective; however, He et al. [2024] focuses on linear regression, while we focus on precision matrix estimation, making the techniques different. Li et al. [2023b], Tian and Feng [2023] studied high-dimensional generalized linear regression. Pathak et al. [2022], Ma et al. [2023], Wang [2023], Ge et al. [2024] addressed covariate shift. Cai and Wei [2021] studied nonparametric classification and Cai et al. [2024a] investigated multi-armed bandit problems. Liu [2023] proposed a unified transfer learning model for high-dimensional linear regression problems. Hanneke and Kpotufe [2019, 2022] studied the transfer learning problem from a learning theory perspective. Lin and Reimherr [2022], Cai et al. [2024b] studied transfer learning for functional data analysis.
The most relevant work to this paper is Li et al. [2023a], which also studied transfer learning for precision matrix estimation. The key difference is the similarity assumption. Li et al. [2023a] assumes that the divergence matrices between the target and source precision matrices are sparse. We assume that most entries of the target precision matrix are shared across source precision matrices, with few different entries. Although the assumption in Li et al. [2023a] is motivated by the KL divergence between Gaussian distributions, ours is a structural assumption, making it applicable beyond Gaussian data and easier to interpret. Consequently, our method differs significantly from that of Li et al. [2023a].
@@
Our approach leverages differential network estimation techniques, which aim to directly estimate the difference between two precision matrices without the need to estimate the individual ones [Zhao et al., 2014, Yuan et al., 2017, Liu et al., 2014, Ma et al., 2021]. Fazayeli and Banerjee [2016] explored this concept in the context of Ising models. Additionally, Zhao et al. [2019, 2022] extended differential network estimation methods to functional data, while Tugnait [2023] broadened its application to multi-attribute data.
@@ 1.2 Organization and Notation
The rest of the paper is organized as following. In Section 2, we introduce the problem setup. In Section 3, we introduce the methodology of the paper. We then describe how to implement our method in practice in Section 4. The theoretical results are developed in Section 5. Besides, we implement extensive simulation experiments in Section 6. Furthermore, in Section 7, we apply our method on two real-world datasets. Finally, we conclude our paper with Section 8. The technical proofs and details about optimization algorithms are provided in the appendix.
@@
[…]
$\mathbb{E}\left[\exp\left(\lambda X_{j}/\sqrt{\Sigma_{jj}}\right)\right]\leq\exp\left(\sigma^{2}\lambda^{2}/2\right)\quad\text{for all }\lambda\in\mathbb{R}\text{ and }1\leq j\leq d,$
where $\Sigma=\textrm{Cov}(X)$.
@@ 2 Problem Setup
Imagine that we observe $n_{0}$ i.i.d. samples $\{\mathbf{x}^{(0)}_{i}\}^{n_{0}}_{i=1}\coloneqq\mathcal{D}_{0}$ from a sub-Gaussian target distribution $\mathcal{P}_{0}$. Each sample $\mathbf{x}^{(0)}_{i}\in\mathbb{R}^{d}$ is assumed to have zero mean and covariance matrix $\Sigma^{(0)}$. Our goal is to estimate the target precision matrix $\Omega^{(0)}=\left\{\Sigma^{(0)}\right\}^{-1}$. Additionally, we have access to $K$ sub-Gaussian source distributions $\{\mathcal{P}_{k}\}^{K}_{k=1}$, each with $n_{k}$ i.i.d. samples $\{\mathbf{x}^{(k)}_{i}\}^{n_{k}}_{i=1}\coloneqq\mathcal{D}_{k}$. For $1\leq k\leq K$, each $\mathbf{x}^{(k)}_{i}\in\mathbb{R}^{d}$ also has zero mean, covariance matrix $\Sigma^{(k)}$, and corresponding precision matrix $\Omega^{(k)}=\left\{\Sigma^{(k)}\right\}^{-1}$. The goal is to leverage samples from both the target and source distributions to accurately estimate $\Omega^{(0)}$.
To facilitate transfer learning, we assume structural similarity between the target and source precision matrices, whereby most entries in $\Omega^{(0)}$ are shared with those in $\Omega^{(k)}$, with relatively few differences. This assumption enables us to efficiently utilize source samples to enhance estimation of the target precision matrix.
Formally, we characterize the relationship between the target and source precision matrices using the following assumption.
@@
where $|\Omega^{\star}|_{0}\leq s$, $|\Gamma^{(k)\star}|_{0}\leq h$, and $|\Gamma^{(k)\star}|_{1}\leq M_{\Gamma}$. The sparsity parameters $s$ and $h$ satisfy $s\gg h$, indicating that the majority of the structure is shared, while unique components are minimal.
See Figure 1 for a visual illustration of Assumption 1.
Assumption 1 is inspired from the assumptions widely used in the differential network estimation literature [Zhao et al., 2014, Yuan et al., 2017, Zhao et al., 2022]. To see this connection, for each $1\leq k\leq K$, define $\Psi^{(k)}=\Omega^{(k)}-\Omega^{(0)}$ to be the differential network between the target $\Omega^{(0)}$ and the source $\Omega^{(k)}$. Two immediate implications of Assumption 1 are
@@
This matches exactly Condition 1 in the paper Zhao et al. [2014].
In addition, Assumption 1 is naturally interpretable within Gaussian graphical models. Suppose we have an undirected graphical model $G=(V,E)$ where nodes represent variables and edges represent conditional dependencies. In this model, an edge exists between nodes $i$ and $j$ if and only if $\Omega^{(k)}_{ij}\neq 0$ [Lauritzen, 1996]. Under Assumption 1, the target precision matrix $\Omega^{(0)}$ and the source matrices $\left\{\Omega^{(k)}\right\}$ share a large subset of edges, with only a small number of unique edges in each source, corresponding to sparse deviations $\Psi^{(k)}$.
@@
Last but not least, we contrast Assumption 1 with the similarity assumptions based on divergence measures, used in Li et al. [2023a]. Li et al. [2023a] assumes that the divergence matrix
$\Upsilon^{(k)}=\Omega^{(0)}\Sigma^{(k)}-I_{d}$ (2)
is sparse. However, our structural approach offers broader applicability beyond Gaussian data, as it does not rely on assumptions specific to Gaussian graphical models and provides a straightforward interpretation of shared structure. See a more detailed comparison in Appendix A.
Next, we discuss estimating $\Omega^{(0)}$ using samples from both target and source distributions under Assumption 1.
@@ 3 Trans-Glasso Algorithm
In this section, we introduce Trans-Glasso, our transfer learning method for precision matrix estimation, which consists of two main steps. First, we initialize estimators of the precision matrices by solving a multi-task learning problem. Second, we refine these estimators using differential network estimation to adjust for structural differences between the target and source matrices.
@@ 3.1 Initialization via Multi-Task Learning
To leverage shared structure across the target and source matrices, we begin by jointly estimating precision matrices for both the target and source distributions. Based on Assumption 1, we employ a multi-task variant of the graphical lasso estimator [Friedman et al., 2008], which we refer to as Trans-MT-Glasso (Transfer Multi-Task Graphical lasso).
Let $\widehat{\Sigma}^{(k)}=\frac{1}{n_{k}}\sum^{n_{k}}_{i=1}\mathbf{x}^{(k)}_{i}\mathbf{x}^{(k)\top}_{i}$ denote the sample covariance matrix for $0\leq k\leq K$. We define $\Theta\coloneqq\left(\Omega,\left\{\Gamma^{(k)}\right\}^{K}_{k=0}\right)$ for the shared component and the sparse unique components. The Trans-MT-Glasso objective is then given by
@@
$\Phi\left(\Theta\right)$ $\coloneqq\left|\Omega\right|_{1}+\sum^{K}_{k=0}\sqrt{\alpha_{k}}\,\left|\Gamma^{(k)}\right|_{1},$ (5)
$\mathcal{C}\left(M_{\mathrm{op}}\right)$$\coloneqq\left\{\Theta=\left(\Omega,\left\{\Gamma^{(k)}\right\}^{K}_{k=0}\right)\,:\,\Omega+\Gamma^{(k)}\succ 0\ \text{and}\ \left\|\Omega+\Gamma^{(k)}\right\|_{2}\leq M_{\mathrm{op}}\ \text{for all}\ 0\leq k\leq K\right\}.$ (6)

### v2

@@ Abstract
Precision matrix estimation is essential in various fields; yet, it is challenging when samples for the target study are limited. Transfer learning can enhance estimation accuracy by leveraging data from related source studies. We propose Trans-Glasso, a two-step transfer learning method for precision matrix estimation. First, we obtain initial estimators using a multi-task learning objective that captures both shared and unique features across studies. Then, we refine these estimators through differential network estimation to adjust for structural differences between the target and source precision matrices. Under the assumption that most entries of the target precision matrix are shared with those of the source matrices, we derive non-asymptotic error bounds and show that Trans-Glasso achieves minimax optimality under certain conditions. Extensive simulations demonstrate Trans-Glasso ’s superior performance compared to baseline methods, particularly in small-sample settings. We further validate Trans-Glasso in applications to gene networks across brain tissues and protein networks for various cancer subtypes, showcasing its effectiveness in biological contexts. Additionally, we derive the minimax optimal rate for differential network estimation, representing the first such guarantee in this area. The Python implementation of Trans-Glasso, along with code to reproduce all experiments in this paper, is publicly available at https://github.com/boxinz17/transglasso-experiments.
@@ 1 Introduction
Estimating the precision matrix, i.e., the inverse covariance matrix, is a fundamental task in statistical analysis and has broad applications, including in portfolio optimization, speech recognition, and genomics (20). The precision matrix is closely tied to Gaussian graphical models: estimating the support of the precision matrix corresponds to uncovering the network structure of conditional dependencies between multivariate normal variables (20). However, estimating a precision matrix accurately is often challenging when the sample size is small compared to the dimension—a typical scenario in high-dimensional settings.
In many applications, the target study has a limited sample size, while data from related studies may be available. Transfer learning (30) provides a promising approach in these scenarios by leveraging information from related source studies to improve estimation accuracy in the target study. For example, in gene expression studies across different tissue types, sample sizes may be small for specific tissues, but data from related tissues can help improve estimates (22). Similarly, protein network studies for different cancer subtypes can benefit from transfer learning, as leveraging data from related subtypes can enhance estimation for a particular subtype with limited data (34).
A critical aspect of transfer learning is establishing similarity between the target and source tasks. Here, we assume that most entries of the target precision matrix are shared with those of the source matrices, with only a few differences. Based on this assumption, we propose Trans-Glasso, a novel two-step transfer learning method for precision matrix estimation. First, we obtain initial estimators through a multi-task learning objective that captures shared and unique dependencies across datasets. Second, we refine these estimators using differential network estimation to adjust for differences between the target and source matrices (50; 45).
We provide a theoretical analysis of Trans-Glasso, deriving non-asymptotic error bounds and establishing that the method achieves minimax optimality in a wide range of parameter settings. Through extensive simulations, we demonstrate that Trans-Glasso outperforms several baseline methods, particularly in scenarios where the target sample size is small. We also apply Trans-Glasso to gene networks across brain tissues and protein networks for various cancer subtypes, showing its practical effectiveness in biological applications. Additionally, as a byproduct of our analysis, we derive the minimax optimal rate for differential network estimation—to our knowledge, the first such guarantee in this area.
@@ 1.1 Related Work
Estimation of sparse precision matrices in a single study is well studied. Common methods include penalized M-estimator (46; 11; 38; 19; 36) and constrained $L_{1}$ minimization (6; 7; 37). There is also extensive literature on multi-task precision matrix estimation, which estimates multiple related but non-identical precision matrices from multiple studies (13; 9; 51). See (40) for a survey. While related to transfer learning, multi-task learning aims to estimate parameters of all studies, whereas transfer learning focuses solely on the target study.
@@
Transfer learning has a long history (30). Recently, interest in transfer learning for statistical problems has grown. Li et al. (21); He et al. (14) studied high-dimensional linear regression. The fused regularizer in (14) is similar to our multi-task objective; however, He et al. (14) focuses on linear regression, while we focus on precision matrix estimation, leading to distinct technical analyses. Li et al. (23); Tian and Feng (39) studied high-dimensional generalized linear regression. Cai and Wei (8) studied nonparametric classification and Cai et al. (4) investigated multi-armed bandit problems. Lin and Reimherr (24); Cai et al. (5) studied transfer learning for functional data analysis. Kpotufe and Martinet (17); Pathak et al. (32); Ma et al. (29); Ge et al. (12); Pathak and Ma (33) addressed covariate shift—a special form of transfer learning—under different statistical models.
The most relevant work to this paper is (22), which also studied transfer learning for precision matrix estimation. The key difference is the similarity assumption. (22) assumes that the divergence matrices between the target and source precision matrices are sparse. We assume that most entries of the target precision matrix are shared across source precision matrices, with few different entries. Although the assumption in (22) is motivated by the KL divergence between Gaussian distributions, ours is a structural assumption, making it applicable beyond Gaussian data and easier to interpret. Consequently, our method differs significantly from that of (22).
@@
Our approach leverages differential network estimation techniques, which aim to directly estimate the difference between two precision matrices without the need to estimate the individual ones (50; 45; 27; 28). Fazayeli and Banerjee (10) explored this concept in the context of Ising models. Additionally, Zhao et al. (48, 47) extended differential network estimation methods to functional data, while Tugnait (42) broadened its application to multi-attribute data.
@@ 1.2 Organization and Notation
The rest of the paper is organized as follows. In Section 2, we introduce the problem setup. In Section 3, we introduce the methodology of the paper. Section 4 describes practical implementation details of our method. The theoretical results are developed in Section 5. We also conduct extensive simulation experiments in Section 6. Furthermore, in Section 7, we apply our method to two real-world datasets. Finally, we conclude our paper with Section 8. The technical proofs and details about optimization algorithms are provided in the appendix. The Python implementation of our method, along with code to reproduce all experiments in this paper, is publicly available at
https://github.com/boxinz17/transglasso-experiments
@@
For a vector $v\in\mathbb{R}^{d}$, we use $\|v\|_{p}$ to denote its $L_{p}$-norm. More specifically, we have $\|v\|_{p}=(\sum^{d}_{i=1}|v_{i}|^{p})^{\frac{1}{p}}$ for $1\leq p\leq\infty$, where $\|v\|_{\infty}=\max_{i}|v_{i}|$. For a matrix $A\in\mathbb{R}^{d\times d}$, we use $|\cdot|$ to denote its elementwise norm and $\|\cdot\|$ to denote its operator norm. For example, $|A|_{1}=\sum^{d}_{i=1}\sum^{d}_{j=1}|A_{ij}|$, $|A|_{0}=\sum^{d}_{i=1}\sum^{d}_{j=1}\mathds{1}\left\{A_{ij}\neq 0\right\}$, $|A|_{\infty}=\max_{1\leq i,j\leq d}|A_{ij}|$; $\|A\|_{1}=\max_{1\leq j\leq d}\sum^{d}_{i=1}|A_{ij}|$, $\|A\|_{\infty}=\max_{1\leq i\leq d}\sum^{d}_{j=1}|A_{ij}|$, and $\|A\|_{2}$ denote the largest singular value of $A$. We use $\|A\|_{\mathrm{F}}=(\sum^{d}_{i=1}\sum^{d}_{j=1}|A_{ij}|^{2})^{1/2}$ to denote the Frobenius norm of $A$. In addition, we use $\langle A,B\rangle=\text{tr}(A^{\top}B)=\sum_{i,j}A_{ij}B_{ij}$ for $A,B\in\mathbb{R}^{d\times d}$ to define the inner product between two matrices. We use $\text{vec}(A)$ to denote the $d^{2}$-vector obtained by stacking the columns of $A$. When $A$ is symmetric, we let $\gamma_{\min}(A)$ and $\gamma_{\max}(A)$ denote its smallest and largest eigenvalue. For $A\in\mathbb{R}^{n_{1}\times n_{2}}$ and $B\in\mathbb{R}^{m_{1}\times m_{2}}$, we let $A\otimes B=[A_{ij}B_{lm}]_{i,j,l,m}\in\mathbb{R}^{n_{1}m_{1}\times n_{2}m_{2}}$ denote the Kronecker product of two matrices. We define $\mathbb{S}^{d\times d}$ as the set of symmetric matrices with dimension $d$. The universal constants may vary from one line to another without further clarification. Finally, following (36), we say a random vector $X\in\mathbb{R}^{d}$ with $\mathbb{E}[X]=0$ and $\Sigma=\textrm{Cov}(X)$ is sub-Gaussian if there exists a constant $\sigma>0$ such that
$\mathbb{E}\left[\exp\left(\lambda X_{j}/\sqrt{\Sigma_{jj}}\right)\right]\leq\exp\left(\sigma^{2}\lambda^{2}/2\right)\quad\text{for all }\lambda\in\mathbb{R}\text{ and }1\leq j\leq d.$
@@ 2 Problem Setup
Imagine that we observe $n_{0}$ i.i.d. samples $\{\mathbf{x}^{(0)}_{i}\}^{n_{0}}_{i=1}\coloneqq\mathcal{D}_{0}$ from a sub-Gaussian target distribution $\mathcal{P}_{0}$. Each sample $\mathbf{x}^{(0)}_{i}\in\mathbb{R}^{d}$ is assumed to have zero mean and covariance matrix $\Sigma^{(0)}$. Our goal is to estimate the target precision matrix $\Omega^{(0)}=(\Sigma^{(0)})^{-1}$. Additionally, we have access to $K$ sub-Gaussian source distributions $\{\mathcal{P}_{k}\}^{K}_{k=1}$, each with $n_{k}$ i.i.d. samples $\{\mathbf{x}^{(k)}_{i}\}^{n_{k}}_{i=1}\coloneqq\mathcal{D}_{k}$. For $1\leq k\leq K$, each $\mathbf{x}^{(k)}_{i}\in\mathbb{R}^{d}$ also has zero mean, covariance matrix $\Sigma^{(k)}$, and corresponding precision matrix $\Omega^{(k)}=(\Sigma^{(k)})^{-1}$. The goal is to leverage samples from both the target and source distributions to accurately estimate $\Omega^{(0)}$.
To facilitate transfer learning, we assume structural similarity between the target and source precision matrices, whereby most entries in $\Omega^{(0)}$ are shared with those in $\Omega^{(k)}$, with relatively few differences. This assumption enables us to efficiently utilize source samples to enhance the estimation of the target precision matrix. Formally, we characterize the relationship between the target and source precision matrices by the following assumption.
@@
where $|\Omega^{\star}|_{0}\leq s$, $|\Gamma^{(k)\star}|_{0}\leq h$, and $|\Gamma^{(k)\star}|_{1}\leq M_{\Gamma}$. The parameters $s$ and $h$ satisfy $s\gg h$, indicating that the majority of the structure is shared, while unique components are minimal.
See Figure 1 for a visual illustration of Assumption 1. Assumption 1 is inspired by the assumptions widely used in the differential network estimation literature (50). To see this connection, for each $1\leq k\leq K$, define $\Psi^{(k)}=\Omega^{(k)}-\Omega^{(0)}$ to be the differential network between the target $\Omega^{(0)}$ and the source $\Omega^{(k)}$. Two immediate implications of Assumption 1 are
@@
This exactly matches Condition 1 in the paper (50).
We also note that the structural similarity in Assumption 1 is conceptually related to the heterogeneity assumption in (1), which models the discrepancy between proxy and true predictive tasks as a sparse bias; both frameworks exploit the idea that differences across tasks are concentrated in a small set of coordinates, thereby enabling efficient transfer.
In addition, Assumption 1 is naturally interpretable within Gaussian graphical models. Suppose we have an undirected graphical model $G=(V,E)$ where nodes represent variables and edges represent conditional dependencies. In this model, an edge exists between nodes $i$ and $j$ if and only if $\Omega^{(k)}_{ij}\neq 0$ (20). Under Assumption 1, the target precision matrix $\Omega^{(0)}$ and the source matrices $\left\{\Omega^{(k)}\right\}$ share a large subset of edges, with only a small number of unique edges in each source, corresponding to sparse deviations $\Psi^{(k)}$.
@@
Finally, we compare Assumption 1 with the similarity assumption based on divergence measures adopted in (22). Specifically, Li et al. (22) assume that the divergence matrix $\Upsilon^{(k)}=\Omega^{(0)}\Sigma^{(k)}-I_{d}$ is sparse, a condition motivated by controlling the KL divergence between two multivariate Gaussian distributions. In contrast, our structural assumption does not rely on Gaussian-specific properties and naturally extends beyond Gaussian graphical models, providing a more intuitive interpretation of shared structure. We emphasize that we do not claim Trans-Glasso consistently outperforms Trans-CLIME, the method proposed in (22), and we leave a detailed theoretical comparison to future work. Further discussion is provided in Appendix A.
@@ 3 Trans-Glasso Algorithm
In this section, we introduce Trans-Glasso, a transfer learning method for precision matrix estimation. It comprises two steps: (1) initializing precision matrix estimators via a multi-task learning problem and (2) refining them with differential network estimation to account for structural differences between target and source matrices.
@@ 3.1 Initialization via Multi-Task Learning
To leverage shared structure across the target and source matrices, we begin by jointly estimating precision matrices for both the target and source distributions. Based on Assumption 1, we employ a multi-task variant of the graphical lasso estimator (11), which we refer to as Trans-MT-Glasso (Transfer Multi-Task Graphical lasso).
Let $\widehat{\Sigma}^{(k)}=\frac{1}{n_{k}}\sum^{n_{k}}_{i=1}\mathbf{x}^{(k)}_{i}\mathbf{x}^{(k)\top}_{i}$ denote the sample covariance matrix for $0\leq k\leq K$. We define $\Theta\coloneqq(\Omega,\left\{\Gamma^{(k)}\right\}^{K}_{k=0})$ for the shared component and the sparse unique components. The Trans-MT-Glasso objective is then given by
@@
$\Phi\left(\Theta\right)$ $\coloneqq\left|\Omega\right|_{1}+\sum^{K}_{k=0}\sqrt{\alpha_{k}}\,\left|\Gamma^{(k)}\right|_{1},\qquad\text{and}$ (4)
$\mathcal{C}\left(M_{\mathrm{op}}\right)$ $\coloneqq\left\{\Theta=\left(\Omega,\left\{\Gamma^{(k)}\right\}^{K}_{k=0}\right)\,:\,\Omega+\Gamma^{(k)}\succ 0\ \text{and}\ \max_{0\leq k\leq K}\left\|\Omega+\Gamma^{(k)}\right\|_{2}\leq M_{\mathrm{op}}\right\}.$ (5)